OpenStudy (dumbsearch2):

A Ferris wheel rotates around in 30 seconds. The maximum height above the ground Is 55 feet, and the minimum height above the ground is 5 feet. Which function models the height as a function of r in seconds?

OpenStudy (dumbsearch2):

@jim_thompson5910

OpenStudy (dumbsearch2):

Here are the possible answer choices: f(t)=25sin(pi/15)t+30 f(t)=25sin2(pi)t+30 f(t)=30sin(pi)/(15)t+25 f(t) = 25sin60(pi)t+30

OpenStudy (anonymous):

Assume at t = 0, height (h) = 5 ft, then: \[h(t) = 5+(25-25*\cos{\frac{\pi}{15} t}) = 30-25*\cos{\frac{\pi}{15} t}\]

OpenStudy (anonymous):

let me see if it is equivalent to some of the answers

OpenStudy (anonymous):

\[h(t)=55-(25-25\sin{\frac{\pi}{15}t}) = 30+25\sin{\frac{\pi}{15}t}\]

OpenStudy (triciaal):